# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ F.1 Determining the Latent Rank from Offline Data
We note that as discussed in 4, we can use the eigenvalues of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to determine the rank of our subspace. We use the version of this arising from pseudo-inverses instead of regularization, just like in the MovieLens experiments. We demonstrate that we can indeed determine that the $d_{K}=18$ by finding the significant eigenvalues of the pseudo-inverse version of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ estimated from the offline dataset of 5000 samples. We show the plots and log plots of these eigenvalues. We also plot the eigenvalues of the completed ratings matrix for comparison. Notice that they match and both fall after $18$ eigenvalues.
@@
Appendix G Additional Figures
@@
Figure 6: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 7: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
@@
Figure 8: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts $1$-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.

### v2

@@ H.1 Determining the Latent Rank from Offline Data
We note that as discussed in 3, we can use the eigenvalues of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ to determine the rank of our subspace. We use the version of this arising from pseudo-inverses instead of regularization, just like in the MovieLens experiments. We demonstrate that we can indeed determine that the $d_{K}=18$ by finding the significant eigenvalues of the pseudo-inverse version of $\overline{\textbf{D}}_{N,1}^{-1}{\overline{\textbf{M}}_{N}}{\overline{\textbf{D}}_{N,2}^{-1}}$ estimated from the offline dataset of 5000 samples. We show the plots and log plots of these eigenvalues. We also plot the eigenvalues of the completed ratings matrix for comparison. Notice that they match and both fall after $18$ eigenvalues.
@@
H.2 Simulation Study
We generate $\textbf{U}_{*}$ with $\textbf{U}_{ij}\text{ i.i.d. }\text{Unif}(0,\frac{2.5}{d_{K}d_{A}})$. We simulate the hidden labels $\bm{\theta}_{n}\sim\mathcal{N}(0,d_{K}^{-1}\textbf{I}_{d_{K}})$, generate feature vectors $\phi(x_{n,h},a_{n,h})\sim\mathcal{N}(0,\textbf{I}_{d_{A}})$ normalized to unit norm, and sample noise $\epsilon_{n,h}\text{ i.i.d. }\mathcal{N}(0,0.5^{2})$. We use SOLD to estimate $\hat{\textbf{U}}$ from the offline dataset $\mathcal{D}_{\operatorname{off}}$, which consists of $5000$ trajectories of length $20$ each. In accordance with the confidence set determined by Li et al., (2010), we choose $\alpha_{1,t}=0.33\sqrt{d_{K}\log(1+10T/d_{K})}$ and $\alpha_{2,t}=0.33\sqrt{d_{A}\log(1+10T/d_{A})}$, and share the LinUCB and ProBALL-UCB hyperparameters by assigning $\alpha_{t}=\alpha_{2,t}$.  [^9] All experiments were run on a single computer with an Intel i9-13900k CPU, 128GB of RAM, and a NVIDIA RTX 3090 GPU, in no more than an hour in total.
@@ H.2 Simulation Study
Figure 5: Comparison of ProBALL-UCB with LinUCB, for different choices of $\tau$ and confidence bound constructions. All variants perform no worse than LinUCB, with martingale Bernstein performing the best. The shaded area depicts 11-standard error confidence intervals over $30$ trials.
H.3 MovieLens
MovieLens (Harper and Konstan,, 2015) is a large-scale movie recommendation dataset comprising $6040$ users and $3883$ movies, where each user may rate one or more movies. Like Hong et al., (2020), we filter the dataset to include only movies rated by at least $200$ users and vice-versa. We factor the sparse rating matrix into user parameters $\bm{\beta}$ and movie features $\Phi$ using the probabilistic matrix factorization algorithm of Mnih and Salakhutdinov, 2007b , using nuclear norm regularization so that the rank of $\bm{\beta}$ is $d_{K}=18$. However, we consider a much higher dimensional problem than Hong et al., (2020) do – we let $d_{A}=200$ so $\bm{\beta}\in\mathbb{R}^{1589\times 200},\Phi\in\mathbb{R}^{200\times 1426}$. At each round for user $i$, the agent chooses between $20$ movies of different genres with features $\Phi_{a_{1}},...,\Phi_{a_{20}}$, and has to recommend the best movie presented to it to maximize the user’s rating of the movie. We generate rewards for recommending movie $j$ to user $i$ by $\bm{\beta}_{i}^{T}\Phi_{j}+\epsilon_{ij},\epsilon_{ij}\text{ i.i.d. }\mathcal{N}(0,0.5)$.
Our hyperparameters are chosen and varied just as in the simulation study. To reproduce the methods of Hong et al., (2020), we cluster the user features into $d_{K}$ clusters using k-means, and provide mUCB and mmUCB with the mean vectors of each cluster as latent models. We initialize ProBALL-UCB with a subspace estimated with an unregularized variant of SOLD, that uses pseudo-inverses instead of inverses, because of difficulties in finding an appropriate regularization parameter for this large, noisy, and high-dimensional dataset. The subspace was estimated from $5000$ trajectories of length $50$ simulated from the reward model and the uniform behavior policy. Note that we assign $\Delta_{\operatorname{off}}$ for $\epsilon$ in mmUCB, as this is their tolerance parameter for model misspecification.
@@ H.3 MovieLens
Figure 6: Comparison of ProBALL-UCB with LinUCB and TS algorithms, for different choices of $\tau$ and confidence bound constructions. All variants perform no worse than LinUCB and outperform the TS algorithms, with martingale Bernstein performing the best. The shaded area depicts 11-standard error confidence intervals over $30$ trials.
Figure 7: Comparison of ProBALL-UCB and ProBALL-TS initialized with SOLD against LinUCB, TS, MixTS, and mmTS, for different choices of $\tau$ and confidence bound constructions. ProBALL-UCB outperforms LinUCB, and ProBALL-TS outperforms MixTS and mmTS. Shaded area depicts 11-standard error confidence intervals over $30$ trials with fresh $\bm{\theta}$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\bm{\theta}$.
H.3.1 UCB Algorithms
Figure 8: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of regret. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 9: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. ProBALL-UCB performs no worse than LinUCB, and outperforms mUCB and mmUCB.
H.3.2 TS Algorithms
Figure 10: Comparison of ProBALL-TS initialized with SOLD against TS, mmTS, and MixTS, for different choices of $\tau$ and confidence bound constructions, in terms of regret. All variants of ProBALL-TS outperform TS, mmTS, and MixTS.
Figure 11: Comparison of ProBALL-TS initialized with SOLD against TS, mmTS, and MixTS, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-TS outperform TS, mmTS, and MixTS. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
H.3.3 Comparison Against Low-Dimensional Ground Truth Subspaces
Figure 12: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB on low-dimensional ground-truth features, for different choices of $\tau$ and confidence bound constructions. When $\tau$ is small enough, all variants of ProBALL-UCB perform no worse than low-dimensional LinUCB, and outperform mUCB and mmUCB, on ground truth features. This showcases the efficacy of SOLD, and demonstrates that we recover subspaces that are just as good as ground-truth. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 13: Comparison of ProBALL-UCB initialized with SOLD against LinUCB, mUCB, and mmUCB on low-dimensional ground-truth features, for different choices of $\tau$ and confidence bound constructions. When $\tau$ is small enough, all variants of ProBALL-UCB perform no worse than low-dimensional LinUCB, and outperform mUCB and mmUCB, on ground truth features. This showcases the efficacy of SOLD, and demonstrates that we recover subspaces that are just as good as ground-truth. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
H.3.4 No Usage of SOLD
Figure 14: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
Figure 15: Comparison of ProBALL-UCB initialized with ground truth subspaces against LinUCB, mUCB, and mmUCB, for different choices of $\tau$ and confidence bound constructions, in terms of rolling average rating over $25$ timesteps. All variants of ProBALL-UCB perform no worse than LinUCB, and outperform mUCB and mmUCB. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$. The confidence intervals on regret thus account for the variation in frequentist regret for changing $\theta$.
H.4 Sample Complexity of SOLD
We perform an empirical study of the sample complexity of SOLD on the MovieLens dataset. To do so, we compare the end-to-end regret at $T=200$ timesteps of both ProBALL-UCB and ProBALL-TS, against LinUCB and Linear Thompson sampling using ground-truth low-dimensional features. When $\tau$ is small enough, we see that the end-to-end regret of both ProBALL-UCB and ProBALL-TS converges to that of LinUCB and Linear Thompson sampling using ground-truth low-dimensional features. This shows that we lose little from needing to estimate the subspace with SOLD when enough offline samples are present.
Figure 16: Subspace estimation error of SOLD against the number of offline samples, in the Frobenius norm. This was performed on the MovieLens dataset. We compare the error of SOLD against the parametric rate of $1/\sqrt{N}$. This shows that the error of SOLD indeed decreases very quickly in practice.
Figure 17: End-to-end regret at $T=200$ timesteps of ProBALL-UCB initialized with SOLD, against the number of offline samples used in fitting SOLD. With a low enough $\tau$, the regret of ProBALL-UCB approaches the regret of LinUCB on ground-truth low-dimensional features, showing that we lose next to nothing from needing to estimate the subspace with SOLD. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$.
Figure 18: End-to-end regret at $T=200$ timesteps of ProBALL-TS initialized with SOLD, against the number of offline samples used in fitting SOLD. With a low enough $\tau$, the regret of ProBALL-TS approaches the regret of TS on ground-truth low-dimensional features, showing that we lose next to nothing from needing to estimate the subspace with SOLD. Shaded area depicts 11-standard error confidence intervals over 30 trials with fresh $\theta$.